If the temperature of a 15.5 mL sample of N₂ gas is increased from 25.0 °C to 75.0 °C as its pressure remains constant, what will be its new volume in milliliters?


1
Expert's answer
2021-06-03T05:28:02-0400

V1 = 15.5 ml .

T1 = 250c = 298 K .

T2 = 750c = 348 K .

V2 = ? .



V1 / T1 = V2 / T2 .


15.5 / 298 = V2 / 348 .


V2 = 15.5 × 348 / 298 ml .


V2 = 18.1006 ml . Answer.